if f(x)=x^3 what is f^-1(1000)
Answer by haileytucki(390)

f(x)=x^(3)_f^(-1)(1000)
To find the inverse of the function, interchange the variables and solve for f^(-1)(x).
x=f^(-1)(x)^(3)
Since f^(-1)(x) is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
f^(-1)(x)^(3)=x
Take the cube root of both sides of the equation to eliminate the exponent on the left-hand side.
f^(-1)(x)=~3:(x)
Evaluate the inverse f^(-1) at x=1000.
f^(-1)(1000)=~3:((1000))
Simplify the function.
f^(-1)(1000)=10